# Localization and interpolation of parabolic $L^{p}$ Neumann problems

### v1

@@ Abstract.
We show a localization estimate for local solutions to the parabolic equation $\partial_{t}u+\operatorname{div}(A\nabla u)=0$ with zero Neumann data, assuming that the $L^{p}$ Neumann problem and $L^{p^{\prime}}$ Dirichlet problem for the adjoint operator are solvable in a Lipschitz cylinder for some $p\in(1,\infty)$. Using this result, we establish the solvability of the Neumann problem in the atomic Hardy space for parabolic operators with bounded, measurable, time-dependent coefficients, and hence obtain the interpolation of solvability of the $L^{p}$ Neumann problem.
@@ 1. Introduction
(1.1) $\|\widetilde{N}(|\nabla u|{\mathds{1}}_{Q})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}},$
for a parabolic cube $Q$ centered at a boundary point such that the Neumann data $\partial^{A}_{\nu}u=0$ in $2Q\cap\partial\Omega$. Here $Lu=0$, where
@@
and the domain $\Omega=\mathcal{O}\times\mathbb{R}$ is a bounded or unbounded Lipschitz cylinder, i.e., $\mathcal{O}=\mathbb{R}^{n}_{+}$ or $\mathcal{O}$ is a bounded or unbounded Lipschitz domain. We assume that $A=[a_{ij}(X,t)]$ is a $n\times n$ bounded matrix satisfying the uniform ellipticity condition for a.e. $X\in\mathcal{O}$, $t\in\mathbb{R}$. That is, there exists positive constants $\lambda$ and $\Lambda$ such that
@@
Let $\Omega=\mathcal{O}\times\mathbb{R}$ where $\Omega$ is either a bounded or unbounded Lipschitz domain in $\mathbb{R}^{n}$. Let $L=-\partial_{t}+\operatorname{div}(A\nabla\cdot)$, and let $L^{*}=\partial_{t}+\operatorname{div}(A^{T}\nabla\cdot)$ be the adjoint operator of $L$. Let $p\in(1,\infty)$. Suppose that $\rm{(}\rm{N}_{p}\rm{)}$L and $\rm{(}\rm{D}_{p^{\prime}}\rm{)}$$^{L^{*}}$ are solvable in $\Omega$ (see Definitions 2.20, 2.21), then for any parabolic cube $Q=Q_{r}(x,t)$ centered at some $(x,t)\in\partial\Omega$, and any local weak solution $u$ to $Lu=0$ in $2Q\cap\Omega$ with zero Neumann data on $2Q\cap\partial\Omega$, we have
$\|\widetilde{N}(|\nabla u|{\mathds{1}}_{Q})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{2Q\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}}.$
@@
For specific operators (such as the heat equation on Lipschitz cylinders, or even more general domains, this result is known (c.f. [4])), but here it is established for operators with no particular assumptions on coefficients beyond ellipticity. Hence this result is more analogous to previously established elliptic theory for the Neumann problem as in [15]. Further, we expect that this interpolation result will be very useful for studying the Neumann problem for classes of specific operators. One such example is our paper in preparation [11], applying the main result here to a subclass of operators of the type considered in [10] and showing optimal $L^{p}$ solvability of the Neumann problem for such PDEs when the Carleson measure condition on the coefficients is assumed to be sufficiently small.
@@ 2. Definitions
A parabolic cube on $\mathbb{R}^{n}\times\mathbb{R}$ centered at $(X,t)$ with sidelength $r$ as
@@ 2.1. Tent spaces
$\mathcal{A}_{1}(w)(q,\tau):=\iint_{\Gamma(q,\tau)}|w(X,t)|\frac{dXdt}{\delta(X,t)^{n+2}},$
@@
$\mathcal{C}_{1}(w)(q,\tau):=\sup_{r>0}\frac{1}{\sigma(B_{r}(q,\tau)\cap\partial\Omega)}\iint_{B_{r}(q,\tau)\cap\Omega}|w(X,t)|\frac{dXdt}{\delta(X,t)}.$
If $w\in L^{p}_{\operatorname{loc}}(\Omega)$, $p\in(0,\infty)$, we define for $c\in(0,1)$ the modified non-tangential maximal function
@@
$\|\widetilde{N}^{c}_{a}(u)\|_{L^{p}(\partial\Omega,\sigma)}\approx\|\widetilde{N}(u)\|_{L^{p}(\partial\Omega,\sigma)},$
@@
where the constant depend on $\Omega$ and $p$.
With those in hand, the space of functions $u$ for which $\|\widetilde{N}^{a,c}(u)\|_{p}$, $\|\widetilde{\mathcal{A}}^{a,c}(u)\|_{p}$ or $\|\widetilde{\mathcal{C}}_{1}^{c}(u)\|_{p}$ are finite doesn’t depend on $a$ or $c$. We call them ($L^{2}$ averaged) tent spaces $\widetilde{T}^{p}_{\infty}(\Omega)$ and $\widetilde{T}^{p}_{1}(\Omega)$ whose norm are
@@
$\|u\|_{\widetilde{T}^{p}_{1}(\Omega)}:=\|\widetilde{\mathcal{A}}_{1}(u)\|_{L^{p}(\partial\Omega,\sigma)}\qquad\text{for }p\in(0,\infty),$
$\|u\|_{\widetilde{T}^{\infty}_{1}(\Omega)}:=\|\widetilde{\mathcal{C}}(u)\|_{L^{\infty}(\partial\Omega,\sigma)}.$
@@ 2.2. Reinforced weak and Energy solutions
Here, $D_{t}^{1/2}$ is the half-order derivative and $H_{t}$ is the Hilbert transform with respect to the $t$ variable, normalized so that $\partial_{t}=D_{t}^{1/2}H_{t}D_{t}^{1/2}$. The space $\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R})$ is the homogeneous Sobolev space of order 1/2 - the completion of $\mathcal{C}_{0}^{\infty}(\mathbb{R})$ in the norm $\|D_{t}^{1/2}(\cdot)\|_{2}$ - and it embeds into the space ${\mathcal{S}}^{\prime}(\mathbb{R})/\mathbb{C}$ of tempered distributions, modulo constants). The local space $\mathcal{H}_{\operatorname{loc}}^{1/2}(\mathbb{R})$ consists of functions $u$ such that $u\phi\in\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R})$ for all $\phi\in\mathcal{C}_{0}^{\infty}(\Omega)$. As shown in [6] the space $\mathcal{H}_{\operatorname{loc}}^{1/2}(\mathbb{R})$ is larger than $\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R})$. For $u\in\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R};L^{2}_{\operatorname{loc}}(\mathcal{O})),$ (2.16) simplifies to
@@
$\iint_{\Omega}A\nabla u\cdot{\nabla\phi}\,\mathrm{d}X\,\mathrm{d}t-\iint_{\Omega}u\cdot{\partial_{t}\phi}\,\mathrm{d}X\,\mathrm{d}t=0.$
@@
Specialising to the case $\Omega=\mathcal{O}\times\mathbb{R}$, where $\mathcal{O}$ is either bounded or unbounded Lipschitz domain, we say that a reinforced weak solution $v\in\dot{\mathsf{E}}_{\operatorname{loc}}(\mathcal{O}\times\mathbb{R})$ belongs to the energy class $\dot{\mathsf{E}}(\mathcal{O}\times\mathbb{R})$ if
@@
As shown in [2] (with a small generalization), functions from $\dot{\mathsf{E}}$ have well defined Dirichlet traces with values in the homogeneous parabolic Sobolev space $\dot{H}{\vphantom{H}}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\mathcal{O}\times\mathbb{R})$. Here, $\dot{H}{\vphantom{H}}^{s}_{\pm\partial_{t}-\Delta_{x}}(\mathbb{R}^{n})$ is defined as the closure of Schwartz functions $v\in{\mathcal{S}}({\mathbb{R}^{n}})$ with Fourier support away from the origin in the norm $\|\mathcal{F}^{-1}((|\xi|^{2}\pm i\tau)^{s}\mathcal{F}v)\|_{2}$. This yields a space of tempered distributions modulo constants in $L_{\operatorname{loc}}^{2}({\mathbb{R}^{n}})$ if $0<s\leq 1/2$. Conversely, any $g\in\dot{H}{\vphantom{H}}^{1/4}_{\partial_{t}-\Delta_{x}}$ can be extended to a function $v\in\dot{\mathsf{E}}$ with trace $v\big|_{\partial\mathbb{R}^{n+1}_{+}}=g$. For this via partition of unity we can define $\dot{H}{\vphantom{H}}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\mathcal{O}\times\mathbb{R})$ for $\mathcal{O}$ Lipschitz.
@@
(2.17) $\iint_{\mathcal{O}\times\mathbb{R}}\left[A\nabla u\cdot{\nabla v}+H_{t}D_{t}^{1/2}u\cdot{D_{t}^{1/2}v}\right]\,\mathrm{d}X\,\mathrm{d}t-\int_{\partial O\times\mathbb{R}}gv\,\mathrm{d}x\,\mathrm{d}t=0,$
@@
Hence, $u+w$ is an energy solution with data $f$. Should there exist another solution $v$, then $a_{\delta}(u+w-v,u+w-v)=0$ and hence by coercivity $\|u+w-v\|_{\dot{\mathsf{E}}}=0$. Thus the two solutions only differ by a constant. It means that Dirichlet problem associated with our parabolic PDE is well-posed in the energy class. Similar arguments allow us to solve the Neumann problem by considering for the datum $g\in\dot{H}{\vphantom{H}}^{-1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$ the solution $u\in\dot{E}$ such that
@@
with zero Dirichlet and Neumann data.
To avoid any boundary terms arising from the pairing of the righthand side with a test function we shall assume that $h$ is a bounded function with bounded support and $F$ is a bounded function with compact support in $\Omega$.
@@
and since $\nabla v\in L^{2}(\Omega)$ we need $\|F\|_{L^{2}(\Omega)}<\infty$, which certainly any bounded compactly supported $F$ satisfies. The above defines a bounded functional on $\dot{\mathsf{E}}_{0}$ for the Dirichlet problem and on $\dot{\mathsf{E}}$ for the Neumann problem and thus we can solve both problems with a righthand side of the form $\operatorname{div}F$.
Considering the righthand side for the term $h$ for the Dirichlet problem is also not problematic but we will need to understand the embedding of the space $\dot{\mathsf{E}}_{0}$ into the $L^{p}$ spaces. It is stated in [2] Remark 3.13 (without proof) that the space $\dot{\mathsf{E}}(\mathbb{R}^{n+1})/\mathbb{C}$ has realization in the Lebesgue space $L^{q}(\mathbb{R}^{n+1})$ for $q=2(n+2)/n$. From this (using an odd extension across the flat boundary) we get that $\dot{\mathsf{E}}_{0}(\mathbb{R}^{n}_{+}\times\mathbb{R})$ embeds into $L^{q}(\mathbb{R}^{n}_{+}\times\mathbb{R})$. Using a partition of unity, the embedding of $\dot{\mathsf{E}}_{0}(\mathcal{O}\times\mathbb{R})$ into $L^{q}(\mathcal{O}\times\mathbb{R})$ follows. Hence it remains to see why the original claim as stated in [2] holds. Taking the same approach as in Tao’s notes [19, Exercise 41] it suffices to understand into which $L^{q}(\mathbb{R}^{n+1})$ belongs the convolution of an $L^{2}$ function with the kernel
@@
It holds that $\rho\sim|\xi|+|\tau|^{1/2}$, and that $f$ is the function $\rho(x,t)^{-1}=\|(x,t)\|^{-1}$. Observe that $\dot{\mathsf{E}}(\mathbb{R}^{n+1})$ norm can be defined by
(2.22) $\|g\|_{\dot{\mathsf{E}}(\mathbb{R}^{n+1})}=\|\mathbb{D}g\|_{L^{2}(\mathbb{R}^{n+1})}=\left\|\|(\xi,\tau)\|\hat{g}(\xi,\tau)\right\|_{L^{2}(\mathbb{R}^{n+1})}.$
This then implies that $g=h*\hat{f}$ for some $L^{2}$ function $h$. Analysis of the kernel $\hat{f}$ above reveals that it scales in $\xi$ as $|\xi|^{-(n+1)}$ but in $\tau$ as $|\tau|^{-(n+1)/2}$. The final conclusion requires understanding of fractional integrals with kernels of mixed homogeneity ( see Chapter 1 for the parabolic norm structure and Theorem 6.1 for the fractional integration result of [12]). This reveals that the resulting convolution belongs to the strong $L^{q}(\mathbb{R}^{n+1})$ for $q=2(n+2)/n$ as desired.

### v2

@@ Abstract.
We show a localization estimate for local solutions to the parabolic equation $\partial_{t}u+\operatorname{div}(A\nabla u)=0$ with zero Neumann data, assuming that the $L^{p}$ Neumann problem and $L^{p^{\prime}}$ Dirichlet problem for the adjoint operator are solvable in a Lipschitz cylinder for some $p\in(1,\infty)$. Using this result, we establish the solvability of the Neumann problem in the atomic Hardy space for parabolic operators with bounded, measurable, time-dependent coefficients, and hence obtain the extrapolation of solvability of the $L^{p}$ Neumann problem.
@@ 1. Introduction
(1.1) $\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{r}})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}},$
for a backward in time parabolic cube $J_{r}$ centered at a boundary point such that the Neumann data $\partial^{A}_{\nu}u=0$ in $J_{2r}\cap\partial\Omega$. Here $Lu=0$, where
@@
and the domain $\Omega=\mathcal{O}\times\mathbb{R}$ is a bounded or unbounded Lipschitz cylinder, i.e., $\mathcal{O}=\mathbb{R}^{n}_{+}$ or $\mathcal{O}$ is a bounded or unbounded Lipschitz domain. We assume that $A=[a_{ij}(X,t)]$ is a $n\times n$ bounded matrix satisfying the uniform ellipticity condition for a.e. $X\in\mathcal{O}$, $t\in\mathbb{R}$. That is, there exist positive constants $\lambda$ and $\Lambda$ such that
@@
Let $\Omega=\mathcal{O}\times\mathbb{R}$ where $\mathcal{O}$ is either a bounded or unbounded Lipschitz domain in $\mathbb{R}^{n}$. Let $L=-\partial_{t}+\operatorname{div}(A\nabla\cdot)$, and let $L^{*}=\partial_{t}+\operatorname{div}(A^{T}\nabla\cdot)$ be the adjoint operator of $L$. Let $p\in(1,\infty)$. Suppose that $\rm{(}\rm{N}_{p}\rm{)}$L and $\rm{(}\rm{D}_{p^{\prime}}\rm{)}$$^{L^{*}}$ are solvable in $\Omega$ (see Definitions 2.22, 2.23). Let $Q_{0}:=Q_{0}^{\prime}\times(T_{0},T_{1})$ be a cube centered at $\partial\Omega$. Then for any $(\bar{x},\bar{t})\in Q_{0}\cap\partial\Omega$, any backward parabolic cube $J_{r}(\bar{x},\bar{t})$ that satisfies $J_{2r}(\bar{x},\bar{t})\subset Q_{0}$, and any weak solution $u$ to $Lu=0$ in $Q_{0}\cap\Omega$ with zero Neumann data on $Q_{0}\cap\partial\Omega$, we have
$\|\widetilde{N}(|\nabla u|{\mathds{1}}_{J_{r}(\bar{x},\bar{t})})\|_{L^{p}(\partial\Omega)}\leq Cr^{(n+1)/p}\left(\mathchoice{{\vbox{\hbox{$\textstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.675pt}}{{\vbox{\hbox{$\scriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-4.02336pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-2.53589pt}}{{\vbox{\hbox{$\scriptscriptstyle\mkern-1.5mu{-}\mkern-7.5mu{-}$}}\kern-1.7709pt}}\!\int\!\!\!\!\!\int_{J_{2r}(\bar{x},\bar{t})\cap\Omega}|\nabla u|^{2}dXdt\right)^{\frac{1}{2}}.$
@@
For specific operators (such as the heat equation on Lipschitz cylinders, or even more general domains), this result is known (cf. [4]), but here it is established for operators with no particular assumptions on coefficients beyond ellipticity. Hence this result is more analogous to previously established elliptic theory for the Neumann problem as in [15]. Further, we expect that this interpolation result will be very useful for studying the Neumann problem for classes of specific operators. One such example is our paper in preparation [11], applying the main result here to a subclass of operators of the type considered in [10] and showing optimal $L^{p}$ solvability of the Neumann problem for such PDEs when the Carleson measure condition on the coefficients is assumed to be sufficiently small.
@@ 2. Definitions
We define a parabolic cube on $\mathbb{R}^{n}\times\mathbb{R}$ centered at $(X,t)$ with sidelength $r$ as
@@
For $(X,t)\in\mathbb{R}^{n}\times\mathbb{R}$, we denote by $J_{r}(X,t)$ the backward in time parabolic cube, that is,
(2.3) $J_{r}(X,t):=Q_{r}^{\prime}(X)\times(t-r^{2},t).$
@@ 2.1. Tent spaces
$\mathcal{A}^{a}_{1}(w)(q,\tau):=\iint_{\Gamma_{a}(q,\tau)}|w(X,t)|\frac{dXdt}{\delta(X,t)^{n+2}},$
@@
$\mathcal{C}_{1}(w)(q,\tau):=\sup_{r>0}\frac{1}{\sigma(B_{r}(q,\tau)\cap\partial\Omega)}\iint_{B_{r}(q,\tau)\cap\Omega}|w(X,t)|\frac{dXdt}{\delta(X,t)},$
where $\sigma=\mathcal{H}^{n+1}_{p}|_{\partial\Omega}$ is the $(n+1)$-dimensional parabolic Hausdorff measure [^1] For a set $E\subset\mathbb{R}^{n+1}$, the $(n+1)$-dimensional parabolic Hausdorff measure is defined as $\mathcal{H}^{n+1}_{p}(E):=\lim_{\epsilon\to 0^{+}}\mathcal{H}^{n+1}_{p,\epsilon}(E)$, where $\mathcal{H}^{n+1}_{p,\epsilon}(E):=\inf\left\{\sum_{i}\operatorname{diam}(E_{i})^{s}:E\subseteq\cup_{i}E_{i},\operatorname{diam}(E_{i})\leq\epsilon\right\},$ and $\operatorname{diam}(E):=\sup_{(X,t),(Y,s)\in E}d_{p}((X,t),(Y,s))$. restricted to $\partial\Omega$. If $w\in L^{p}_{\operatorname{loc}}(\Omega)$222with respect to the $(n+1)$-dimensional Lebesgue measure, $p\in(0,\infty)$, we define for $c\in(0,1)$ the modified non-tangential maximal function
@@
$\|\widetilde{N}^{c}_{a,2}(u)\|_{L^{p}(\partial\Omega,\sigma)}\approx\|\widetilde{N}(u)\|_{L^{p}(\partial\Omega,\sigma)},$
@@
where the constants depend on $\Omega$ and $p$.
With those in hand, the space of functions $u$ for which $\|\widetilde{N}^{c}_{a,2}(u)\|_{p}$, $\|\widetilde{\mathcal{A}}^{a,c}_{1,2}(u)\|_{p}$ or $\|\widetilde{\mathcal{C}}_{1,2}^{c}(u)\|_{p}$ are finite doesn’t depend on $a$ or $c$. We call them ($L^{2}$ averaged) tent spaces $\widetilde{T}^{p}_{\infty}(\Omega)$ and $\widetilde{T}^{p}_{1}(\Omega)$ whose norms are
@@
$\|u\|_{\widetilde{T}^{p}_{1}(\Omega)}:=\|\widetilde{\mathcal{A}}(u)\|_{L^{p}(\partial\Omega,\sigma)},\qquad\text{for }p\in(0,\infty),\quad\|u\|_{\widetilde{T}^{\infty}_{1}(\Omega)}:=\|\widetilde{\mathcal{C}}(u)\|_{L^{\infty}(\partial\Omega,\sigma)}.$
@@ 2.2. Reinforced weak and Energy solutions
Here, $D_{t}^{1/2}$ is the half-order derivative and $H_{t}$ is the Hilbert transform with respect to the $t$ variable, normalized so that $\partial_{t}=D_{t}^{1/2}H_{t}D_{t}^{1/2}$. The space $\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R})$ is the homogeneous Sobolev space of order 1/2 - the completion of $\mathcal{C}_{0}^{\infty}(\mathbb{R})$ in the norm $\|D_{t}^{1/2}(\cdot)\|_{2}$ - and it embeds into the space ${\mathcal{S}}^{\prime}(\mathbb{R})/\mathbb{C}$ of tempered distributions, modulo constants. The local space $\mathcal{H}_{\operatorname{loc}}^{1/2}(\mathbb{R})$ consists of functions $u$ such that $u\phi\in\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R})$ for all $\phi\in\mathcal{C}_{0}^{\infty}(\Omega)$. As shown in [6] the space $\mathcal{H}_{\operatorname{loc}}^{1/2}(\mathbb{R})$ is larger than $\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R})$. For $u\in\dot{H}{\vphantom{H}}^{1/2}(\mathbb{R};L^{2}_{\operatorname{loc}}(\mathcal{O})),$ (2.17) simplifies to
@@
(2.18) $\iint_{\Omega}A\nabla u\cdot{\nabla\phi}\,\mathrm{d}X\,\mathrm{d}t-\iint_{\Omega}u\cdot{\partial_{t}\phi}\,\mathrm{d}X\,\mathrm{d}t=0.$
@@
Specialising to the case $\Omega=\mathcal{O}\times\mathbb{R}$, where $\mathcal{O}$ is either a bounded or an unbounded Lipschitz domain, we say that a reinforced weak solution $v\in\dot{\mathsf{E}}_{\operatorname{loc}}(\mathcal{O}\times\mathbb{R})$ belongs to the energy class $\dot{\mathsf{E}}(\mathcal{O}\times\mathbb{R})$ if
@@
As shown in [2] (with a small generalization), functions from $\dot{\mathsf{E}}$ have well defined Dirichlet traces with values in the homogeneous parabolic Sobolev space $\dot{H}{\vphantom{H}}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\mathcal{O}\times\mathbb{R})$. Here, $\dot{H}{\vphantom{H}}^{s}_{\pm\partial_{t}-\Delta_{x}}(\mathbb{R}^{n})$ is defined as the closure of Schwartz functions $v\in{\mathcal{S}}({\mathbb{R}^{n}})$ with Fourier support away from the origin in the norm $\|\mathcal{F}^{-1}((|\xi|^{2}\pm i\tau)^{s}\mathcal{F}v)\|_{2}$. This yields a space of tempered distributions modulo constants in $L_{\operatorname{loc}}^{2}({\mathbb{R}^{n}})$ if $0<s\leq 1/2$. Conversely, any $g\in\dot{H}{\vphantom{H}}^{1/4}_{\partial_{t}-\Delta_{x}}$ can be extended to a function $v\in\dot{\mathsf{E}}$ with trace $v\big|_{\partial\mathbb{R}^{n+1}_{+}}=g$. From this via partition of unity we can define $\dot{H}{\vphantom{H}}^{1/4}_{\partial_{t}-\Delta_{x}}(\partial\mathcal{O}\times\mathbb{R})$ for $\mathcal{O}$ Lipschitz.
@@
(2.19) $\iint_{\mathcal{O}\times\mathbb{R}}\left[A\nabla u\cdot{\nabla v}+H_{t}D_{t}^{1/2}u\cdot{D_{t}^{1/2}v}\right]\,\mathrm{d}X\,\mathrm{d}t-\int_{\partial\mathcal{O}\times\mathbb{R}}gv\,\mathrm{d}x\,\mathrm{d}t=0,$
@@
Hence, $u+w$ is an energy solution with data $f$. Should there exist another solution $v$, then $a_{\delta}(u+w-v,u+w-v)=0$ and hence by coercivity $\|u+w-v\|_{\dot{\mathsf{E}}}=0$. Thus the two solutions only differ by a constant. Thus the Dirichlet problem associated with our parabolic PDE is well-posed in the energy class. Similar arguments allow us to solve the Neumann problem by considering for the datum $g\in\dot{H}{\vphantom{H}}^{-1/4}_{\partial_{t}-\Delta_{x}}(\partial\Omega)$ the solution $u\in\dot{E}$ such that
@@
with zero Dirichlet or Neumann data.
To avoid any boundary terms arising from the pairing of the right-hand side with a test function we shall assume that $h$ is a bounded function with bounded support and $F$ is a bounded function with compact support in $\Omega$.
@@
and since $\nabla v\in L^{2}(\Omega)$ we need $\|F\|_{L^{2}(\Omega)}<\infty$, which certainly any bounded compactly supported $F$ satisfies. The above defines a bounded functional on $\dot{\mathsf{E}}_{0}$ for the Dirichlet problem and on $\dot{\mathsf{E}}$ for the Neumann problem and thus we can solve both problems with a right-hand side of the form $\operatorname{div}F$.
Considering the right-hand side for the term $h$ for the Dirichlet problem is also not problematic but we will need to understand the embedding of the space $\dot{\mathsf{E}}_{0}$ into the $L^{p}$ spaces. It is stated in [2] Remark 3.13 (without proof) that the space $\dot{\mathsf{E}}(\mathbb{R}^{n+1})/\mathbb{C}$ has realization in the Lebesgue space $L^{q}(\mathbb{R}^{n+1})$ for $q=2(n+2)/n$. From this (using an odd extension across the flat boundary) we get that $\dot{\mathsf{E}}_{0}(\mathbb{R}^{n}_{+}\times\mathbb{R})$ embeds into $L^{q}(\mathbb{R}^{n}_{+}\times\mathbb{R})$. Using a partition of unity, the embedding of $\dot{\mathsf{E}}_{0}(\mathcal{O}\times\mathbb{R})$ into $L^{q}(\mathcal{O}\times\mathbb{R})$ follows. Hence it remains to see why the original claim as stated in [2] holds. Taking the same approach as in Tao’s notes [19, Exercise 41] it suffices to understand into which $L^{q}(\mathbb{R}^{n+1})$ belongs the convolution of an $L^{2}$ function with the kernel
@@
It holds that $\rho\sim|\xi|+|\tau|^{1/2}$, and that $f$ is the function $\rho(x,t)^{-1}=\|(x,t)\|^{-1}$. Observe that the $\dot{\mathsf{E}}(\mathbb{R}^{n+1})$ norm can be defined by
$\|g\|_{\dot{\mathsf{E}}(\mathbb{R}^{n+1})}=\|\mathbb{D}g\|_{L^{2}(\mathbb{R}^{n+1})}=\left\|\|(\xi,\tau)\|\hat{g}(\xi,\tau)\right\|_{L^{2}(\mathbb{R}^{n+1})},$
where
$(\mathbb{D}f)\,\widehat{\,}\,(\xi,\tau):=\|(\xi,\tau)\|\widehat{f}(\xi,\tau).$
This then implies that $g=h*\hat{f}$ for some $L^{2}$ function $h$. Analysis of the kernel $\hat{f}$ above reveals that it scales in $\xi$ as $|\xi|^{-(n+1)}$ but in $\tau$ as $|\tau|^{-(n+1)/2}$. The final conclusion requires understanding of fractional integrals with kernels of mixed homogeneity (see Chapter 1 for the parabolic norm structure and Theorem 6.1 for the fractional integration result of [12]). This reveals that the resulting convolution belongs to the strong $L^{q}(\mathbb{R}^{n+1})$ space for $q=2(n+2)/n$, as desired.